OpenStudy (anonymous):

What is f(x)=-3(x+2)^2+4 in standard form? How do I rewrite it ?

Vocaloid (vocaloid):

just expand (x+2)^2 = (x+2)*(x+2) = ____

OpenStudy (anonymous):

I'm really stupid please explain??

Vocaloid (vocaloid):

use foil to multiply (x+2)*(x+2)